Need an explanation please.

Prove it mathematically and in words why a launch angle of 45 degrees will provide the max range of a projectile.
1
Expert's answer
2019-10-23T09:49:10-0400

The horizontal range of a projectile is "d = \frac{v^2 sin 2\theta}{g}".

For a fixed "v" and "g", in order to maximize "d", we need to maximize "\sin 2 \theta". Since "\sin x" attains its maximum value at "x = \frac{\pi}{2}", "\sin 2 \theta" is maximized at "\theta = \frac{\pi}{4}".